The tens digit of a two-digit number exceeds its units digit by 4. The number exceeds twice the number obtained by reversing the digits of the original number by 10. What is the original number?
Answer by stanbon(75887)

The tens digit of a two-digit number exceeds its units digit by 4. The number exceeds twice the number obtained by reversing the digits of the original number by 10. What is the original number?
----------------------
Let the number be 10t+u
---
Equations:
t = u +4
10t+u = 2(10u+t)+10
--------------------------
Simplify the 2nd equation:
8t-19u = 10
t = u + 4
----------------
Substitute for "t" and solve for "u":
8(u+4) - 19u = 10
-11u + 32 = 10
-11u = -22
u = 2
----
Since t = u+4, t = 6
========================
Original Number: 10t+u = 62
==============================
Cheers,
Stan H.